# Symmetrical Flow Matching: Unified Image Generation, Segmentation, and Classification with Score-Based Generative Models

### v1

@@ Introduction
Figure 5: Non-curated samples generated by the model trained on CelebAMask-HQ (top) and COCO-stuff (bottom). The top row shows the semantic mask used to condition the model. The bottom row shows the samples after 25 integration steps with the Euler ODE solver.
Classification
@@
Table 3 presents the classification performance comparison. With a single inference step, SymmFlow achieves accuracy comparable to the Diffusion Classifier while being significantly more efficient. By increasing the number of steps to just 25, which is 100 times fewer than required by the Diffusion Classifier, SymmFlow convincingly outperforms it on CIFAR-10. This highlights the efficiency of the approach, reducing inference time without sacrificing accuracy. Additionally, the model currently uses a simple conditioning strategy based on grayscale intensities, suggesting further improvements can be achieved with more advanced conditioning mechanisms. In the Appendix, Figures 7 and 8 present non-curated samples generated by SymmFlow on MNIST and CIFAR-10, illustrating the model’s ability to produce diverse and visually coherent outputs. These findings indicate that SymmFlow continues to be a robust image generator, enabling efficient conditional control.
@@
Impact of Inference Steps
Table 4 evaluates the effect of the number of inference steps on semantic image generation performance across both CelebAMask-HQ and COCO-Stuff. On both datasets, FID and LPIPS scores steadily decrease as the number of steps increases, indicating improvements in image fidelity and structural coherence. In both datasets, the LPIPS trend is particularly revealing: at low step counts, the high LPIPS values do not primarily reflect meaningful diversity, but rather poor visual quality and weak alignment with the conditioning masks. As the number of steps increases, the model produces more coherent and semantically accurate outputs, leading to a decrease in LPIPS. Importantly, the final LPIPS values remain relatively high, but this now reflects genuine variability across samples rather than noise or misalignment, indicating that the model preserves diversity while improving structural fidelity. Figures 9 and 10 in the Appendix illustrate the qualitative effects of using different numbers of steps for sampling.
@@
Dataset Metric 1 2 5 10 20 25
MNIST Acc. $\uparrow$ 99.3 99.4 99.5 99.4 99.5 99.6
CIFAR-10 Acc. $\uparrow$ 88.2 52.3 63.5 74.9 89.4 90.6
For the MNIST data, this effect is less pronounced. Although evolving towards a Gaussian distribution causes class boundaries to blur, the simplicity of digit shapes allows for correct classification across different step counts. These results suggest that although single-step inference can be effective, datasets with more complex semantics benefit from a sufficient number of steps to preserve meaningful class information.
Limitations & Future Work
@@
One limitation of the current approach is the overall model size. Although SymmFlow uses far fewer inference steps than typical diffusion models, it relies on a large pre-trained Stable Diffusion U-Net backbone, making the total model size substantial. Reducing this computational burden, for example, by distilling the model into a one-step or otherwise more efficient variant, would be a valuable direction for future work. Additionally, fine-tuning the VAE decoder to better align with semantic masks shows promise for improving segmentation accuracy on fine-grained details, such as small or occluded regions.
@@
Future work includes extending classification evaluation from the current proof-of-concept stage to datasets such as Food-101 (Bossard, Guillaumin, and Van Gool 2014), FGVC-Aircraft (Maji et al. 2013), Oxford-IIIT Pets (Parkhi et al. 2012), Flowers102 (Nilsback and Zisserman 2008), ImageNet-1K (Deng et al. 2009), and ObjectNet (Barbu et al. 2019), while also refining the semantic label encoding strategy to improve conditioning. Furthermore, evaluating the model performance on depth estimation tasks, similar to DepthFM, would further demonstrate its versatility. Beyond segmentation, the bi-directional formulation of the model presents opportunities for applications such as image editing.
Conclusions
@@
This work introduces Symmetrical Flow Matching, a unified framework that models segmentation, classification, and image generation as opposing flows within a single architecture. Leveraging a bi-directional formulation, SymmFlow enables efficient semantic reasoning while preserving the flexibility required for high-fidelity generation. Unlike prior approaches that impose rigid one-to-one mappings, it supports diverse conditioning strategies, including pixel-level and image-level supervision. Experimental results show that SymmFlow achieves state-of-the-art performance in semantic image synthesis with only 25 inference steps, competitive segmentation accuracy despite operating in a low-resolution latent space, and promising results on classification. These findings demonstrate that large-scale flow-based models can effectively bridge discriminative and generative tasks within a single system. Future work will further explore its classification potential, and extend the framework to structured prediction problems such as depth estimation and semantic image editing.
Appendix A Appendix/supplemental material
@@
The supplementary material is organized as follows: Appendix B covers the datasets used in this work. Appendix C describes the implementation details of the employed models and the compute resources required for training and evaluating the models. Appendix D contains additional qualitative results of sampling and segmentation.
Appendix B Datasets
@@
This section provides details on the datasets used for classification and semantic segmentation and generation. For classification, we use MNIST and CIFAR-10, while for segmentation, we evaluate on CelebAMask-HQ and COCO-Stuff. The datasets are summarized in Table 7.
@@
MNIST (Deng 2012) 60,000 10,000 10
CIFAR-10 (Krizhevsky, Nair, and Hinton 2010) 50,000 10,000 10
CelebAMask-HQ (Lee et al. 2020) 24,183 2,824 19
COCO-stuff (Caesar, Uijlings, and Ferrari 2018) 118,287 5,000 171
The classification datasets, MNIST and CIFAR-10, consist of low-resolution images (32$\times$32), which provide a controlled setting for evaluating fundamental classification capabilities. In contrast, the segmentation datasets, CelebAMask-HQ and COCO-Stuff, contain higher-resolution images (512$\times$512), allowing for a more detailed assessment of semantic segmentation performance.
@@
Appendix C Implementation Details
@@
For pixel-space implementations on MNIST and CIFAR-10, we use the U-Net architecture proposed by Dhariwal et al. (Dhariwal and Nichol 2021), which has been widely adopted for diffusion-based generative modeling. The architecture consists of a series of residual blocks, self-attention layers, and group normalization, enabling effective denoising and feature extraction across multiple resolutions. Table 8 presents the hyperparameters used for training these two models. To ensure reproducibility, we will also make the code publicly available.
@@
For the latent-space implementations, we leverage the pre-trained Variational Autoencoder (VAE) from Stable Diffusion, which efficiently compresses high-dimensional image data into a lower-dimensional latent space while preserving perceptual quality. The VAE can be downloaded from HuggingFace  [^3] https://huggingface.co/stabilityai/sd-vae-ft-mse. The U-Net and pretrained weights correspond to the ones used on Stable Diffusion 2.1 and available on HuggingFace  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1. After loading the weights, the number of input channels in the first layer and the number of output channels in the last layer are doubled. These models were trained with mixed precision, while supporting multi-GPU training. This was made possible through the Accelerate library. Table 9 presents the hyperparameters used for training these two models.
@@
Appendix D Additional Qualitative Results
Figure 7 contains conditional samples generated by the model trained on MNIST. Figure 8 shows conditional samples from the model trained on CIFAR-10. Figures 9 and 10 demonstrate the effect of sampling steps in the sample quality for CelebAMask-HQ and COCO-stuff, respectively.

### v2

@@ 4.6 Implementation Details
Figure 4: Non-curated samples generated by the model trained on CelebAMask-HQ (left) and COCO-stuff (right). The top row shows the semantic mask used to condition the model. The bottom row shows the samples after 25 integration steps with the Euler ODE solver.
5 Results & Discussion
5.1 Semantic Segmentation and Generation
Table 2 reports semantic segmentation and image synthesis results across benchmarks. SymmFlow consistently outperforms prior models in the synthesis task, achieving the lowest FID scores across datasets. The visualizations in Figure 4 confirm that the model produces mask-consistent high-fidelity samples on CelebAMask-HQ and COCO-Stuff, capturing structural details with strong adherence to the conditioning masks. Although LPIPS is commonly used to assess perceptual diversity, it can be misleading in isolation, as higher LPIPS values may result from poor image quality rather than true variability. Conversely, low LPIPS may indicate mode collapse or data leakage rather than accurate diversity. Therefore, interpreting LPIPS jointly with FID is essential. SymmFlow achieves the best trade-off between image quality and diversity, reflecting strong generative capability without sacrificing semantic alignment.
In the segmentation task, SymmFlow achieves competitive performance compared to specialized segmentation baselines, particularly on COCO-Stuff. As shown in Figure 5, the model demonstrates semantic understanding beyond the ground-truth annotations; for instance, correctly identifying a laptop absent from the label map. However, performance is limited by the low-resolution latent representation ($\times64$$\times$4), which hampers fine-grained accuracy. This is especially evident for small-area classes such as earrings or partially occluded features like eyebrows and ears, where segmentation quality deteriorates due to insufficient spatial detail. Despite this, the global segmentation structure remains coherent. SemFlow, while also achieving strong performance on the segmentation task, falls short on semantic image synthesis. Quantitative results indicate limited visual fidelity, underscoring the limitations of the architecture in jointly modeling segmentation and generation without further design or training adjustments.
SymmFlow achieves this performance using only 25 function evaluations, whereas most diffusion-based methods require hundreds of denoising steps, resulting in significantly improved inference efficiency, as shown in Appendix C.
5.2 Classification
@@
Table 3 presents the classification performance comparison. With a single inference step, SymmFlow achieves accuracy comparable to the Diffusion Classifier while being significantly more efficient. By increasing the number of steps to just 25, which is 100 times fewer than required by the Diffusion Classifier, SymmFlow convincingly outperforms it on CIFAR-10. This highlights the efficiency of the approach, reducing inference time without sacrificing accuracy. Additionally, the model currently uses a simple conditioning strategy based on grayscale intensities, suggesting further improvements can be achieved with more advanced conditioning mechanisms. In Appendix G, Figures 6 and 7 present non-curated samples generated by SymmFlow on MNIST and CIFAR-10, illustrating the model’s ability to produce diverse and visually coherent outputs. These findings indicate that SymmFlow continues to be a robust image generator, enabling efficient conditional control.
@@
5.3 Evaluation of Inference Steps
Table 4 evaluates the effect of the number of inference steps on semantic image generation performance across both CelebAMask-HQ and COCO-Stuff. On both datasets, FID and LPIPS scores steadily decrease as the number of steps increases, indicating improvements in image fidelity and structural coherence. In both datasets, the LPIPS trend is particularly revealing: at low step counts, the high LPIPS values do not primarily reflect meaningful diversity, but rather poor visual quality and weak alignment with the conditioning masks. As the number of steps increases, the model produces more coherent and semantically accurate outputs, leading to a decrease in LPIPS. Importantly, the final LPIPS values remain relatively high, but this now reflects genuine variability across samples rather than noise or misalignment, indicating that the model preserves diversity while improving structural fidelity. Appendix E analyzes the impact of the dequantization factors, while Appendix G illustrates the qualitative effects of different numbers of sampling steps.
@@ 5.3 Evaluation of Inference Steps
The impact of the inference steps in classification is covered in Appendix F, and summarized in Table 11.
6 Limitations & Future Work
@@
One limitation of the current approach is the overall model size. Although SymmFlow uses far fewer inference steps than typical diffusion models, it relies on a large pre-trained Stable Diffusion U-Net backbone. Reducing the computational burden by distilling the model into a one-step variant is a natural next step. Moreover, a one-step formulation could facilitate the integration of more expressive conditioning mechanisms, enabling effective text-based control and the use of more powerful architectures such as MMDiT (esser2024scaling). Additionally, fine-tuning the VAE decoder to better align with semantic masks could improve segmentation accuracy on fine-grained details, such as small or occluded regions.
@@
Future work includes extending classification evaluation from the current proof-of-concept stage to datasets such as Food-101 (bossard2014food), ImageNet-1K (deng2009imagenet), and ObjectNet (barbu2019objectnet), while also refining the semantic label encoding strategy to improve conditioning. Furthermore, evaluating the model performance on depth estimation tasks, similar to DepthFM, would further demonstrate its versatility. Beyond segmentation, the bi-directionality of the model presents opportunities for applications such as image editing.
7 Conclusions
@@
This work introduces Symmetrical Flow Matching, a unified framework that models segmentation, classification, and image generation as opposing flows within a single architecture. Leveraging a bi-directional formulation, SymmFlow enables efficient semantic reasoning while preserving the flexibility required for high-fidelity generation. Unlike prior approaches that impose rigid one-to-one mappings, it supports various conditioning strategies, including pixel- and image-level supervision. Experimental results show that SymmFlow achieves state-of-the-art performance in semantic image synthesis with only 25 inference steps, competitive segmentation accuracy despite operating in a low-resolution latent space, and promising classification results. These results show that flow-based generative models can simultaneously support generation and discrimination within a unified framework. Future work will further explore its classification potential and extend the framework to problems such as depth estimation and semantic image editing.
Acknowledgments
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-13640. This research was funded by the European Xecs Eureka TASTI Project.
Appendix/supplemental material
@@ Acknowledgments
The supplementary material is organized as follows: Appendix A covers the datasets used in this work. Appendix B describes the implementation details of the employed models and the computational resources required for training and evaluating the models. Appendix C offers details on model image generation latency. Appendix D provides additional information about how the semantic image synthesis metrics are computed. Appendix E analyses the impact of the dequantization term $\beta$ in the image generation performance. Appendix F evaluates the impact of inference steps on classification performance. Appendix G contains additional qualitative results of sampling.
Appendix A Datasets
@@
This section provides details on the datasets used for classification and semantic segmentation and generation. For classification, we use MNIST (deng2012mnist) and CIFAR-10 (krizhevsky2010cifar), while for segmentation, we evaluate on CelebAMask-HQ (lee2020maskgan) and COCO-Stuff (caesar2018coco). The datasets are summarized in Table 6.
@@
MNIST 60,000 10,000 10
CIFAR-10 50,000 10,000 10
CelebAMask-HQ 24,183 2,824 19
COCO-stuff 118,287 5,000 171
The classification datasets, MNIST and CIFAR-10, consist of low-resolution images ($\times32$), which provide a controlled setting for evaluating fundamental classification capabilities. In contrast, the segmentation datasets, CelebAMask-HQ and COCO-Stuff, contain higher-resolution images ($\times512$), allowing for a more detailed assessment of semantic segmentation performance.
@@
Appendix B Implementation Details
@@
For pixel-space implementations on MNIST and CIFAR-10, we use the U-Net architecture proposed by Dhariwal et al. (dhariwal2021diffusion), which has been widely adopted for diffusion-based generative modeling. The architecture consists of a series of residual blocks, self-attention layers, and group normalization, enabling effective denoising and feature extraction across multiple resolutions. Table 7 presents the hyperparameters used for training these two models. To ensure reproducibility, we will also make the code publicly available.
@@
For latent-space implementations, we use the pre-trained Variational Autoencoder (VAE) from Stable Diffusion, which efficiently compresses high-dimensional image data into a lower-dimensional latent space while preserving perceptual quality. The VAE can be downloaded from HuggingFace  [^3] https://huggingface.co/stabilityai/sd-vae-ft-mse. The U-Net and pretrained weights correspond to those used in Stable Diffusion 2.1 and are available on HuggingFace  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1. After loading the weights, the number of input channels in the first layer and the number of output channels in the last layer are doubled. These models were trained with mixed precision while supporting multi-GPU training. This was made possible through the Accelerate library. Table 8 presents the training hyperparameters for both models.
@@
Appendix C Image Generation Latency
This experiment measures the average time required to generate a single image, including both the forward pass through the VAE and the generator. Latency was evaluated for ControlNet, SemFlow, and SymmFlow, as these models achieve competitive performance with a practical number of inference steps. Other baselines were excluded from this comparison due to either excessively high computational cost or insufficient generation quality, as observed in the one-step variants. The reported results reflect the mean generation time per image under identical hardware conditions, on a system featuring an NVIDIA A100-SXM4 GPU with 40GB of VRAM, an Intel Xeon Platinum 8360Y CPU, and 512GB of RAM.
The results in Table 9 show that SymmFlow achieves a substantial reduction in generation time compared to ControlNet. While SymmFlow and SemFlow exhibit nearly identical inference times, the former delivers significantly higher visual quality and consistency.
Table 9: Image generation latency, measured as the average time required to generate a single image.
Model Latency (ms)
ControlNet (zhang2023adding) 5,483.8
SemFlow (wang2024semflow) 2,140.1
SymmFlow (Proposed) 2,156.9
Appendix D Metrics
For the quantitative evaluation of semantic image synthesis, 50,000 images are generated from the available validation masks, which are used to compute perceptual metrics. The FID is evaluated by comparing the generated samples with the real images from the dataset, assessing overall visual realism and distribution alignment. The LPIPS metric is computed against the validation set by comparing each synthesized image to its corresponding ground-truth image derived from the same mask. For the CIFAR-10 dataset, 50,000 images are sampled instead, evenly distributed across the ten classes (5,000 per class) to ensure balanced evaluation.
Appendix E Dequantization Impact on Generation
This experiment evaluates the impact of different dequantization factors $\beta$ on semantic image generation performance for the CelebAMask-HQ and COCO-Stuff datasets. CelebAMask-HQ was trained with a dequantization factor of $\beta=10.0$, while COCO-Stuff used $\beta=6.0$. The results in Table 10 demonstrate that a moderate increase in $\beta$ at inference time can yield improved generation quality, as reflected by lower FID values. This observation suggests that adjusting the dequantization strength can help balance sample diversity and reconstruction fidelity. A more systematic investigation of $\beta$ during training is necessary to better characterize its influence on the generative process and the resulting image quality.
Table 10: Semantic image generation performance for different values of $\beta$ on CelebAMask-HQ and COCO-Stuff. These results were obtained after 25 integration steps with the Euler ODE solver.
Dataset Metric 4.0 6.0 8.0 10.0 12.0
[…]
Appendix F Inference Steps on Classification
Dataset Metric 1 2 5 10 20 25
MNIST Acc. $\uparrow$ 99.3 99.4 99.5 99.4 99.5 99.6
CIFAR-10 Acc. $\uparrow$ 88.2 52.3 63.5 74.9 89.4 90.6
For the MNIST data, this effect is less pronounced. Although evolving towards a Gaussian distribution causes class boundaries to blur, the simplicity of digit shapes allows for correct classification across different step counts. These results suggest that, although single-step inference can be effective, datasets with more complex semantics benefit from a sufficient number of steps to preserve meaningful class information.
Appendix G Additional Qualitative Results
Figure 6 contains conditional samples generated by the model trained on MNIST. Figure 7 shows conditional samples from the model trained on CIFAR-10. Figures 8 and 9 demonstrate the effect of sampling steps on the sample quality for CelebAMask-HQ and COCO-stuff, respectively. Furthermore, Figure 10 showcases the diversity of the samples produced by SymmFlow, when conditioned on the same semantic mask.
@@ Appendix G Additional Qualitative Results
Figure 10: Non-curated images generated by the model trained on CelebAMask-HQ (left) and COCO-stuff (right). The top row shows the semantic mask used to condition the model. The other rows show the generated samples after 25 integration steps with the Euler ODE solver.